OpenStudy (anonymous):

Which equation matches the statement “The product of 3 and the difference of a number and 10 is 15?"

OpenStudy (anonymous):

3 + (x – 10) = 15 3x – 10 = 15 3 (x – 10) = 15 3 = 15 are answers

OpenStudy (anonymous):

possible answers

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

3 (x – 10) = 15

OpenStudy (anonymous):

3(x-10) = 15

OpenStudy (anonymous):

thanks